OpenStudy (anonymous):

Use law of sins A=42 degrees a= 29 inches b = 21 inches Find B?

OpenStudy (anonymous):

The law of sines states that a/(sin(A)) = b/(sin(B)) = c/(sin(C)) Can you figure it out from there? :)

OpenStudy (anonymous):

I found B =29 degrees

OpenStudy (anonymous):

but what is B' ?

OpenStudy (anonymous):

when B = 29 degrees what is B' ?

OpenStudy (anonymous):

29/sin(42) = 21/B Solve for B

OpenStudy (anonymous):

You already found the anser B=28.98 ~29

OpenStudy (anonymous):

yeah but

OpenStudy (anonymous):

what is B with the apostrophe means ?

OpenStudy (anonymous):

 This is the conventional oblique triangle diagram. You have solved for B. The lowercase letters represent the sides and the uppercase are the angles.